OpenStudy (anonymous):

Solve graphically: 3y= 5-x and 2x = y+3. Write down the co-ordinates of the point of intersection of the two lines.

OpenStudy (anonymous):

y=-1/3x+5/3 and y=2x-3 and graph

OpenStudy (anonymous):

they cross at (2,1) you can see the solution by looking

OpenStudy (anonymous):

\[3y= 5-x\]\[2x = y+3\] \[y=\frac{5}{3}-\frac{1}{3}x\]\[y=2x-3\] \[\frac{5}{3}-\frac{1}{3}x=2x-3\]\[\frac{5-x}{3}=2x-3\]\[5-x=3(2x-3)\]\[5-x=6x-9\]\[7x=14\]\[x=2\] \[y=2\times2-3\]\[y=1\] So, the point of intersection of the two lines is (2,1).